determine the integral of function (ln(4x))^3 /x dx


1
Expert's answer
2021-08-12T15:31:17-0400

Let us determine the integral


"\int\frac{(\ln(4x))^3}{x} dx
=\int(\ln(4x))^3 d(\ln(4x))
=\frac{(\ln(4x))^4}{4}+C."